OpenStudy (anonymous):

Solve the equation. Show your work. 163 = 4x

OpenStudy (anonymous):

its 16^3=4x

OpenStudy (anonymous):

take cube of 16 then divide both sides by 4 that will be ans

OpenStudy (anonymous):

x=16^3/4

OpenStudy (anonymous):

so would it be 1024 divied by four

OpenStudy (anonymous):

or would 1024 be my answer

OpenStudy (anonymous):

16^3=?@jbarber

OpenStudy (anonymous):

16^3=16*16*16=4096

OpenStudy (anonymous):

then 4096/4=1024 am i right or wrong

OpenStudy (anonymous):

divide by 4 =1024 its your ans

OpenStudy (anonymous):

ok thats whati got thank you can u help me with another uestion plz